The denominator fraction is 8 more than its numenator. On subtracting one from numenator and addding 2 to the denominator, the fraction become 2/5. Find the fraction

Grade:11

1 Answers

Arun
5 years ago
Dear Bipin
 
Let the fraction is x/y
Now
y = x+ 8
(x-1) /(y +2) = 2/5
5x - 5 = 2y +4
5x -2y = 9
Solve both equations
5x - 2x - 16 = 9
3x = 25
x = 25/3
Hence
Find y and then fraction will be x/y